OpenStudy (anonymous):

WHAT IS 1 4/5 DIVIDED BY 3/14

OpenStudy (anonymous):

square root of 5 to the 4 power divided 3/14 use wolfram alpha.com to help check it

OpenStudy (anonymous):

HUH I DONT GET IT SORRY?

OpenStudy (anonymous):

\[\sqrt[5]{1}^{4}\] divide by 3/14

OpenStudy (anonymous):

I GOT 27/70

OpenStudy (anonymous):

is it correct?

OpenStudy (anonymous):

idk

OpenStudy (anonymous):

i got 196

OpenStudy (anonymous):

how?

OpenStudy (anonymous):

go to wolframalpha.com to check

OpenStudy (anonymous):

no sorry

OpenStudy (phi):

\[ 1 \frac{4}{5} \div \frac{3}{14} \] change the first number to an improper fraction 1 and 4/5 is 9/5 (1 is 5/5 and add in 4/5) instead of dividing, flip the 2nd fraction and multiply \[ \frac{9}{5}\cdot \frac{14}{3} \] you can simplify by dividing top and bottom by 3: \[ \frac{\cancel{9}3}{5}\cdot \frac{14}{\cancel{3}1} = \frac{42}{5}\] or \[ 8\frac{2}{5} \]